# Quantum Mikheyev–Smirnov–Wolfenstein effect

### v1

@@ Abstract
We study neutrino oscillations in background matter within the quantum field theory formalism where neutrino mass eigenstates are virtual particles. In this case, neutrino mass eigenstates are mixed owing to the interaction with matter. Assuming that neutrinos are Majorana particles, we find the exact propagators for massive neutrinos accounting for the interaction with matter by solving the analog of the Dyson equation. These propagators are used to calculate the transition probability which coincides with the prediction of the Mikheyev–Smirnov–Wolfenstein effect. Finally, we analyze the approximations made in our analysis.
@@ 1 Introduction
We mention one of these approaches based on the Quantum Field Theory (QFT), which was proposed in Refs. [8, 9]. It consists in the consideration of neutrinos, propagating from a source to a detector, as virtual particles. Some basic details are outlined shortly in Sec. 2. Then, this formalism was extensively studied, e.g., in Refs. [10, 11, 12, 13, 14].
@@
In the present work, we study neutrino flavor oscillations in matter based on the QFT approach, i.e. we consider neutrino mass eigenstates as virtual particles. Despite of the difficulties mentioned above, we manage to account for the interaction of the mass eigenstates with matter exactly. Thus, based on reasonable approximations, we find the exact propagators of the mass eigenstates in matter. Eventually, we reproduce the MSW effect using our results.
This work is organized in the following way. We start in Sec. 2 with a brief description of the formalism for neutrino flavor oscillations within QFT. Then, in Sec. 3, we remind how mixed neutrinos interact with background matter. In Sec. 4, considering two flavor case, we derive the equations for the exact propagators of massive neutrinos in background matter. The developed formalism is adapted for Majorana neutrinos in Sec. 5. The exact propagators for massive Weyl neutrinos in matter are found in Sec. 6. These propagators are used in Sec. 7 to find the transition probability for neutrino flavor oscillations in matter within the QFT approach. Finally, we conclude in Sec. 8. The propagators of massive Weyl neutrinos accounting for the diagonal interaction with matter are derived in Appendix A. In Appendix B, we compute the specific 3D Fourier transforms of the propagators obtained in Appendix A.
@@ 2 Quantum field theory formalism for neutrino flavor oscillations
$S=-\frac{1}{2}\left(\sqrt{2}G_{F}\right)^{2}\int\mathrm{d}^{4}x\mathrm{d}^{4}y\left\langle l_{\alpha}\left|T\left\{j_{\mu}^{\dagger}(x)J^{\mu}(x)j^{\nu}(y)J_{\nu}^{\dagger}(y)\right\}\right|l_{\beta}\right\rangle,$ (2.3)
@@
$\hat{l}_{\lambda}(y)\left|l_{\beta}\right\rangle=\frac{\delta_{\lambda\beta}}{\sqrt{2VE_{\beta}}}e^{-\mathrm{i}p_{\beta}y}u(p_{\beta})\left|0\right\rangle,\quad\left\langle l_{\alpha}\right|\hat{\bar{l}}_{\lambda}(x)=\frac{\delta_{\lambda\beta}}{\sqrt{2VE_{\alpha}}}e^{\mathrm{i}p_{\alpha}x}\bar{u}(p_{\alpha})\left\langle 0\right|$ (2.4)
@@
Based on the Dirac matrices in the chiral representation, one transforms the factor in the integrand in Eq. (2.7),
@@
The remaining integrals in Eq. (2.7) can be calculated using the new variables $t=x_{0}-y_{0}$ and $T=(x_{0}+y_{0})/2$. The final result has the form,
@@
$\mathcal{M}_{\beta\to\alpha}=\sum_{ab}U_{\alpha a}U_{\beta b}^{*}\kappa_{-}^{\dagger}(p_{\alpha})\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\Sigma_{ab}(E,\mathbf{q})e^{\mathrm{i}\mathbf{qL}}\kappa_{-}(p_{\beta}).$ (2.10)
@@ 3 Mixed neutrinos in background matter
For simplicity, we consider the system of two mixed active flavor neutrinos $\nu_{\lambda}$, e.g., $\lambda=e,\mu$, in presence of background matter. This matter is supposed to be electroneutral, nonmoving, and unpolarized. The neutrino interaction with matter is treated in the forward scattering approximation. In this case, the Lagrangian for this system has the form,
@@
where $g\equiv g_{12}\neq 0$. One can see in Eq. (3.5) that the mass eigenstates are converted into each other in background matter.
@@ 4 Exact propagators of massive neutrinos in matter
One can see in Eqs. (3.5) that it is problematic to solve the equation for a Green function even for two neutrinos. Indeed, one should consider the multiplet $\Psi=(\psi_{1},\psi_{2})$ and has to invert the $8\times 8$ matrix in Dirac indices while looking for a Green function. Nevertheless, we can tackle the problem of finding the propagators of mass eigenstates perturbatively and try to sum all the terms in a perturbation series.
@@
In Eqs. (4.2) and (4.3), $S_{1,2}=S_{1,2}(p)$ are the 4D Fourier images of the propagators $S_{1,2}(x)$. The same notice refers to $\Sigma_{ab}$. Note that the treatment of neutrino oscillations in matter analogous to that in Eqs. (4.2) and (4.3) was made in Ref. [22]. The analytical expressions in Eqs. (4.2) and (4.3) are illustrated in Fig. 2.
Figure 2: Feynman diagrams corresponding to Eqs. (4.2) and (4.3). 4D Fourier images of the bare propagators $-\mathrm{i}S_{a}$ are depicted by thin solid lines, whereas thick solid lines correspond to the dressed propagators $-\mathrm{i}\Sigma_{ab}$ [23].
The sum of the series in Eqs. (4.2) and (4.3) can be represented in the form of the Dyson equation for an exact Green function,
@@
The solutions of Eq. (4.4) is quite nontrivial especially for Dirac neutrinos.
One can notice that Eq. (4.5) implies the inversion of the effective potential $G$. However, it contains the projection $(1-\gamma^{5})/2$ for Dirac neutrinos. It means that $G^{-1}$ does not exist. Nevertheless, the developed formalism is well defined for Majorana particles. This result is in agreement with the statement (see, e.g., Refs. [24, 25]) that neutrino mass eigenstates are likely to be Majorana particles.
@@ 5 Reformulation of the formalism for Majorana mass eigenstates
The analog of Eq. (3.5) for Weyl neutrinos is (see also Appendix A)
@@
The relation between $(g_{1,2},g)$ and $V_{e,\mu}$ is the same as for Dirac neutrinos. The main results of Sec. 4, e.g., Eqs. (4.4) and (4.5), remain valid for Weyl neutrinos as well. However, $G=g$ is the scalar quantity.
@@ 6 Exact propagators of massive Weyl neutrinos
In this section, we solve Eqs. 4.4 and 4.5 for Majorana neutrinos.
@@
where both $\mathcal{A}_{a}$ and $\mathcal{B}_{a}$ contain four terms. Then, after tedious but straightforward calculations, one finds the solution of Eqs. (4.4) and (4.5),
@@
Further analysis of Eq. (6.2) is quite difficult since $\mathcal{A}_{a}$ and $\mathcal{B}_{a}$ have the complicated structure. However, as shown in Appendix B, the main contribution to the matrix element involving the 3D Fourier transform in Eq. (B.1) comes from $I_{1}$ and $J_{1}$ terms. That is why, we keep only one term in $\mathcal{A}_{a}$ and $\mathcal{B}_{a}$, i.e.
@@
To proceed, we assume that the interaction with matter is weak. In this approximation, we keep the terms up to $\sim g^{2}$ in Eq. (6.2). Using Eq. (6.3), we rewrite Eq. (6.2) in the form,
@@ 7 Transition probability for oscillations in matter
where the exact propagators $\Sigma_{ab}(p_{0},\mathbf{p})$ are given in Eq. (6.4). To get the final form of the matrix element in Eq. (7.1) one should put $p_{0}=E$, calculate the 3D Fourier transform, as well as average the result over the incoming and outgoing left-handed charged leptons.
@@
Analogously to Appendix B, we assume that the neutrino propagation distance is large in deriving of Eq. (7.9). If the incoming and outgoing leptons are ultrarelativistic and their momenta are almost along the same direction, the quantity $\kappa_{-}^{\dagger}(p_{\mu})\left[1-(\bm{\sigma}\hat{L})\right]\kappa_{-}(p_{e})$ is just a nonzero constant factor.
@@
$g_{1}=V_{e}\cos^{2}\theta+V_{\mu}\sin^{2}\theta,\quad g_{1}=V_{e}\sin^{2}\theta+V_{\mu}\cos^{2}\theta,\quad g=(V_{e}-V_{\mu})\sin\theta\cos\theta,$ (7.10)
@@
$P_{e\to\mu}\propto$ $\frac{\left(\frac{\Delta m^{2}}{4E}\sin 2\theta\right)^{2}}{\left(\frac{\Delta m^{2}}{4E}\sin 2\theta\right)^{2}+\left(\frac{\Delta m^{2}}{4E}\cos 2\theta-\frac{G_{\mathrm{F}}n_{e}}{\sqrt{2}}\right)^{2}}$
@@
which coincides with the prediction of the MSW effect.
@@ 8 Conclusion
In Sec. 4, we have derived the equations for the exact propagators, which account for the neutrino interaction with matter. These equations are analogous to the Dyson equations. Moreover, we have established that the developed formalism is not applicable for the Dirac mass eigenstates. For this reason, in Sec. 5, we have reformulated the basic concepts of our approach in terms of Weyl spinors, which are the representation of Majorana neutrinos. The exact propagators for this kind of neutrinos have been found in Sec. 6.
In Sec. 7, we have applied of findings to a particular case of $\nu_{e}\to\nu_{\mu}$ oscillations. We have derived the matrix element and the corresponding transition probability. One can see in Eq. (7.11) that it reproduces the MSW effect.
@@
Despite it is difficult to get something beyond Eq. (7.11) which was originally derived within the quantum mechanical approach, it is instructive to apply the exact formalism of QFT for neutrino flavor oscillations mainly from the methodological point of view. It allows one to analyze the validity of approximations made to obtain Eq. (7.11). The main assumptions used in our work are (i) that neutrinos are ultrarelativistic particles; and (ii) the weakness of the matter interaction.
The assumption (i) was used first in Sec. 5, where it allowed us to exclude the involvement of antineutrinos in the series in Eqs. (4.2) and (4.3). Without this assumption, the summation of the diagrams in Fig. 1 would be quite nontrivial. Second, the assumption (i) was utilized in Appendix B in the computation of the integrals $I_{i}$ and $J_{i}$. It allowed us to show that the contribution of $I_{1}$ and $J_{1}$ is the leading one. Thus, we could reduce the number of terms in $S_{a}$; cf. Eq. (6.3). Third, the assumption (i) was used in the calculation of the 3D Fourier transform in Eqs. (7.6)-(7.8) resulting in the matrix element in Eq. (7.9). Thus, we can see that the ultrarelativity of neutrinos is crucial for Eq. (7.11).
The assumption (ii) was used to simplify the propagators and jump from Eq. (6.2) to Eq. (6.4). In fact, it allowed us to get the $g^{2}$-term under the square root in Eq. (7.3). In principle, one can study corrections to Eq. (7.11) for neutrinos propagating in extremely dense matter.

### v2

# Quantum field theory treatment of neutrino flavor oscillations in matter

@@ Abstract
We study neutrino oscillations in background matter within the quantum field theory formalism where neutrino mass eigenstates are virtual particles. In this case, neutrino mass eigenstates are mixed owing to the interaction with matter. Assuming that neutrinos are Majorana particles, we find the exact propagators for massive neutrinos accounting for the interaction with matter by solving the analog of the Dyson equation. These propagators are used to calculate the transition probability which coincides with the prediction of the standard quantum mechanical treatment of neutrino flavor oscillations in uniform matter. Finally, we analyze the approximations made in our analysis.
@@ 1 Introduction
We mention one of these approaches based on the Quantum Field Theory (QFT), which was proposed in Refs. [8, 9]. It consists in the consideration of neutrinos, propagating from a source to a detector, as virtual particles. Some basic details are outlined shortly in Sec. 2. Then, this formalism was extensively studied, e.g., in Refs. [10, 11, 12, 13, 14, 15].
@@
In the present work, we study neutrino flavor oscillations in matter based on the QFT approach, i.e. we consider neutrino mass eigenstates as virtual particles. Inspite of the difficulties mentioned above, we manage to account for the interaction of the mass eigenstates with matter exactly. Thus, based on reasonable approximations, we find the exact propagators of the mass eigenstates in matter. Eventually, using our results, we reproduce the transition probability for neutrino flavor oscillations in uniform matter obtained within the quantum mechanical approach.
This work is organized in the following way. We start in Sec. 2 with a brief description of the formalism for neutrino flavor oscillations within QFT. Then, in Sec. 3, we remind how mixed neutrinos interact with background matter. In Sec. 4, considering two flavor case, we derive the equations for the exact propagators of massive neutrinos in background matter. The developed formalism is adapted for Majorana neutrinos in Sec. 5. The exact propagators for massive Weyl neutrinos in matter are found in Sec. 6. These propagators are used in Sec. 7 to find the transition probability for neutrino flavor oscillations in matter within the QFT approach. The detailed comparison of QFT and quantum mechanical approaches is provided in Sec. 7.1. Finally, we conclude in Sec. 8. The propagators of massive Weyl neutrinos accounting for the diagonal interaction with matter are derived in Appendix A. In Appendix B, we compute the specific 3D Fourier transforms of the propagators obtained in Appendix A.
@@ 2 Quantum field theory formalism for neutrino flavor oscillations
$S=-\frac{1}{2}\left(\sqrt{2}G_{\mathrm{F}}\right)^{2}\int\mathrm{d}^{4}x\mathrm{d}^{4}y\left\langle l_{\alpha}\left|T\left\{j_{\mu}^{\dagger}(x)J^{\mu}(x)j^{\nu}(y)J_{\nu}^{\dagger}(y)\right\}\right|l_{\beta}\right\rangle,$ (2.3)
@@
$\hat{l}_{\lambda}(y)\left|l_{\beta}\right\rangle=\frac{\delta_{\lambda\beta}}{\sqrt{2VE_{\beta}}}e^{-\mathrm{i}p_{\beta}y}u(p_{\beta})\left|0\right\rangle,\quad\left\langle l_{\alpha}\right|\hat{\bar{l}}_{\lambda}(x)=\frac{\delta_{\lambda\beta}}{\sqrt{2VE_{\alpha}}}e^{\mathrm{i}p_{\alpha}x}\bar{u}(p_{\alpha})\left\langle 0\right|,$ (2.4)
@@
Based on the Dirac matrices in the chiral representation, one transforms the factor in the integrand in Eq. (2),
@@
The remaining integrals in Eq. (2) can be calculated using the new variables $t=x_{0}-y_{0}$ and $T=(x_{0}+y_{0})/2$. The final result has the form,
@@
$\mathcal{M}_{\beta\to\alpha}=\sum_{ab}U_{\alpha a}U_{\beta b}^{*}\kappa_{-}^{\dagger}(p_{\alpha})\left(\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\Sigma_{ab}(E,\mathbf{q})e^{\mathrm{i}\mathbf{qL}}\right)\kappa_{-}(p_{\beta}).$ (2.10)
@@ 3 Mixed neutrinos in background matter
For simplicity, we consider the system of two mixed active flavor neutrinos $\nu_{\lambda}$, e.g., $\lambda=e,\mu$, in the presence of background matter. This matter is supposed to be electroneutral, nonmoving, and unpolarized. The neutrino interaction with matter is treated in the forward scattering approximation. In this case, the Lagrangian for this system has the form,
@@
where $g\equiv g_{12}\neq 0$. One can see in Eq. (3) that the mass eigenstates are converted into each other in background matter.
@@ 4 Exact propagators of massive neutrinos in matter
One can see in Eqs. (3) that it is problematic to solve the equation for a Green function even for two neutrinos. Indeed, one should consider the multiplet $\Psi=(\psi_{1},\psi_{2})$ and has to invert the $8\times 8$ matrix in Dirac indices while looking for a Green function. Nevertheless, we can tackle the problem of finding the propagators of mass eigenstates perturbatively and try to sum all the terms in a perturbation series.
@@
In Eqs. (4) and (4), both $S_{1,2}=S_{1,2}(p)$ and $\Sigma_{ab}$ are the 4D Fourier images of the propagators $S_{1,2}=S_{1,2}(x)$ and $\Sigma_{ab}(x)$ respectively. Note that the treatment of neutrino oscillations in matter analogous to that in Eqs. (4) and (4) was made in Ref. [23]. The analytical expressions in Eqs. (4) and (4) are illustrated in Fig. 2.
Figure 2: Feynman diagrams corresponding to Eqs. (4) and (4). 4D Fourier images of the bare propagators $-\mathrm{i}S_{a}$ are depicted by thin solid lines, whereas thick solid lines correspond to the dressed propagators $-\mathrm{i}\Sigma_{ab}$ [24].
The sum of the series in Eqs. (4) and (4) can be represented in the form of the Dyson equation for an exact Green function,
@@
The diagonal propagators $S_{1,2}$, which are in Eq. (4), were found in Ref. [21] for the case of Dirac neutrinos. They are $4\times 4$ matrices with the quite nontrivial Dirac matrices structure. Thus, despite the solutions of Eq. (4), in principle, exist for Dirac neutrinos, finding them may be rather complicated in practice. One can also notice that Eq. (4) implies the inversion of the effective potential $G$. However, it contains the projection $(1-\gamma^{5})/2$ for Dirac neutrinos. It means that $G^{-1}$ does not exist in this situation. Therefore, the treatment of neutrino flavor oscillations, proposed in our work, cannot be used for Dirac neutrinos.
Nevertheless, the developed formalism is well defined for Majorana particles. It was found, e.g., in Refs. [25, 26] that the diagonalization of the neutrino mass matrix of the general form, which includes both left and right Majorana terms, as well as the Dirac term, gives one certain number of Majorana mass eigenstates. Moreover, the most natural explanation of the small masses of active neutrinos implies that mass eigenstates are Majorana particles (see, e.g., Ref. [27]). It is the heuristic argument for considering Majorana neutrinos in our work.
It is known (see, e.g., Ref. [28]) that a Dirac mass eigenstate can be treated as two Majorana neutrinos with equal masses. However, such Dirac neutrinos cannot be studied in frames of our approach since it requires the consideration of more than two types of particles. As one can see in Eq. (3), it is essential in our method that one deals with only two massive degrees of freedom. We also analyze this approximation in Sec. 8.
@@ 5 Reformulation of the formalism for Majorana mass eigenstates
The analog of Eq. (3) for Weyl neutrinos is (see also Appendix A)
@@
The relation between $(g_{1,2},g)$ and $V_{e,\mu}$ is the same as for Dirac neutrinos. The main results of Sec. 4, e.g., Eqs. (4) and (4), remain valid for Weyl neutrinos as well. However, $G=g$ is the scalar quantity.
@@ 6 Exact propagators of massive Weyl neutrinos
In this section, we solve Eqs. (4) and (4) for Majorana neutrinos.
@@
where both $\mathcal{A}_{a}$ and $\mathcal{B}_{a}$ contain four terms. Then, after tedious but straightforward calculations, one finds the solution of Eqs. (4) and (4),
@@
Further analysis of Eq. (6) is quite difficult since $\mathcal{A}_{a}$ and $\mathcal{B}_{a}$ have the complicated structure. However, as shown in Appendix B, the main contribution to the matrix element involving the 3D Fourier transform in Eq. (B.1) comes from $I_{1}$ and $J_{1}$ terms. That is why, we keep only one term in $\mathcal{A}_{a}$ and $\mathcal{B}_{a}$, i.e.
@@
To proceed, we assume that the interaction with matter is weak. In this approximation, we keep the terms up to $\sim g^{2}$ in Eq. (6). Using Eq. (6.3), we rewrite Eq. (6) in the form,
@@ 7 Transition probability for oscillations in matter
where the exact propagators $\Sigma_{ab}(p_{0},\mathbf{p})$ are given in Eq. (6). To get the final form of the matrix element in Eq. (7.1) one should put $p_{0}=E$, calculate the 3D Fourier transform, as well as average the result over the incoming and outgoing left-handed charged leptons.
@@
Analogously to Appendix B, we assume that the neutrino propagation distance is large in deriving of Eq. (7). If the incoming and outgoing leptons are ultrarelativistic and their momenta are almost along the same direction, the quantity $\kappa_{-}^{\dagger}(p_{\mu})\left[1-(\bm{\sigma}\hat{L})\right]\kappa_{-}(p_{e})$ is just a nonzero constant factor.
@@
$g_{1}=V_{e}\cos^{2}\theta+V_{\mu}\sin^{2}\theta,\quad g_{2}=V_{e}\sin^{2}\theta+V_{\mu}\cos^{2}\theta,\quad g=(V_{e}-V_{\mu})\sin\theta\cos\theta,$ (7.10)
@@
$P_{e\to\mu}=$ $\frac{\left(\frac{\Delta m^{2}}{4E}\sin 2\theta\right)^{2}}{\left(\frac{\Delta m^{2}}{4E}\sin 2\theta\right)^{2}+\left(\frac{\Delta m^{2}}{4E}\cos 2\theta-\frac{G_{\mathrm{F}}n_{e}}{\sqrt{2}}\right)^{2}}$
@@
which coincides with the prediction of the standard quantum mechanical treatment for neutrino flavor oscillations in uniform background matter.
7.1 Comparison with the quantum mechanical approach
Originally, in Ref. [3], the neutrino propagation and oscillations in matter were studied within the quantum mechanical approach. First, we discuss the case of a uniform matter. Considering two flavor eigenstates, $\nu_{f}^{\mathrm{T}}=(\nu_{e},\nu_{\mu})$, one gets that their evolution obeys the effective Schrödinger equation [29],
$\mathrm{i}\frac{\mathrm{d}\nu_{f}}{\mathrm{d}x}=H_{\mathrm{eff}}\nu_{f},\quad H_{\mathrm{eff}}=\left(\begin{array}[]{cc}-\frac{\Delta m^{2}}{4E}\cos 2\theta+\sqrt{2}G_{\mathrm{F}}n_{e}&\frac{\Delta m^{2}}{4E}\sin 2\theta\\ \frac{\Delta m^{2}}{4E}\sin 2\theta&\frac{\Delta m^{2}}{4E}\cos 2\theta\end{array}\right),$ (7.12)
where we assume that the beam of ultrarelativistic neutrinos propagates in space, i.e. $x\approx t$.
If $n_{e}$ in Eq. (7.12) does not depend on coordinates, the transition probability for the process $\nu_{e}\to\nu_{\mu}$, based on Eq. (7.12), coincides with that in Eq. (7) when $x=L$. Now, we assume that the matter density changes adiabatically from a great value to zero. In this case, we can split the trajectory into thin layers, with the density in any of them being approximately constant. If a neutrino beam travels through such a matter, it will reach the layer with the density $n_{e}^{(\mathrm{res})}=\tfrac{\Delta m^{2}\cos 2\theta}{2\sqrt{2}G_{\mathrm{F}}E}$, where the transition probability significantly increases. This phenomenon is called the resonant amplification of neutrino oscillations in matter, or the MSW effect.
The consideration of coordinate dependent matter density is problematic with the QFT approach studied here. Firstly, the propagators, obtained in Appendix A, have quite complicated form if $g_{a}=g_{a}(\mathbf{x})$. Second, the QFT formalism, as it was developed in Refs. [8, 9], implies the derivation of probabilities rather than an evolution equation, like in Eq. (7.12). It is impossible to get the analytical transition probability for the arbitrarily dependent matter density. Nevertheless, we can consider the expression in Eq. (7) as the probability within a thin layer with approximately constant density provided the adiabaticity condition is fulfilled [29].
The relativistic quantum mechanics approach, where neutrinos are treated as first quantized fields, was developed in Refs. [18, 19, 20]. It allows one to rederive the effective Schrödinger equation for neutrino oscillations in various external fields using the exact solutions of the Dirac equation accounting for these external fields. Hence, in that approach, one can consider time and coordinate dependent external fields (see, e.g., Refs. [30, 31]). However, as mentioned in Sec. 1, the relativistic quantum mechanics description of neutrino oscillations has a shortcoming of a quite broad initial neutrino wavepacket.
@@ 8 Conclusion
In Sec. 4, we have derived the equations for the exact propagators, which account for the neutrino interaction with matter. These equations are analogous to the Dyson equations. Moreover, we have established that the developed formalism is not applicable for the Dirac mass eigenstates. For this reason, in Sec. 5, we have reformulated the basic concepts of our approach in terms of Weyl spinors, which are the representation of Majorana neutrinos. The exact propagators for this kind of neutrinos have been found in Sec. 6. Despite sizable experimental efforts are made (see, e.g., Ref. [32]), the issue whether neutrinos are Dirac or Majorana particles is still open. Nevertheless, Majorana neutrinos are preferable from the theoretical point of view since they provide a natural explanation of small masses of active neutrinos [27].
In Sec. 7, we have applied our findings to a particular case of $\nu_{e}\to\nu_{\mu}$ oscillations. We have derived the matrix element and the corresponding transition probability. One can see in Eq. (7) that it reproduces the transition probability for flavor oscillations in matter. We have compared our QFT based approach with the standard quantum mechanical description of neutrino oscillations in Sec. 7.1.
@@
Despite it is difficult to get something beyond Eq. (7) which was originally derived within the quantum mechanical approach, it is instructive to apply the exact formalism of QFT for neutrino flavor oscillations mainly from the methodological point of view. It allows one to analyze the validity of approximations made to obtain Eq. (7). The main assumptions used in our work are (i) that neutrinos are ultrarelativistic particles; and (ii) the weakness of the matter interaction.
The assumption (i) was used first in Sec. 5, where it allowed us to exclude the involvement of antineutrinos in the series in Eqs. (4) and (4). Without this assumption, the summation of the diagrams in Fig. 1 would be quite nontrivial. Second, the assumption (i) was utilized in Appendix B in the computation of the integrals $I_{i}$ and $J_{i}$. It allowed us to show that the contribution of $I_{1}$ and $J_{1}$ is the leading one. Thus, we could reduce the number of terms in $S_{a}$; cf. Eq. (6.3). Third, the assumption (i) was used in the calculation of the 3D Fourier transform in Eqs. (7.6)-(7.8) resulting in the matrix element in Eq. (7). Thus, we can see that the ultrarelativity of neutrinos is crucial for Eq. (7).
The assumption (ii) was used to simplify the propagators and jump from Eq. (6) to Eq. (6). In fact, it allowed us to get the $g^{2}$-term under the square root in Eq. (7.3). In principle, one can study corrections to Eq. (7) for neutrinos propagating in extremely dense matter.
There are a couple of additional important assumptions used in our work. The original formulation of the QFT method in Refs. [8, 9] implied that one gets the probabilities rather than the evolution equation. That is why, we assume that the matter density is constant. The discussion of the possibility to take into account a nonuniform matter has been provided in Sec. 7.1.
We have also studied the system of two neutrinos. It allowed us to derive the analytical transition probability in Eq. (7). One can notice in Fig. 2 that the matter interaction induces the transitions like $1\to 2$ or $2\to 1$. It makes possible to sum the series in Eqs. (4) and (4). If more mass eigenstates were considered, it would be difficult to derive general Eqs. (4) and (4) since the Feynman graphs would include greater number of transitions.
An additional reason why oscillations of Dirac neutrinos cannot be studied within the described approach results from the aforementioned argument. Indeed, in case of Dirac neutrinos, one has to consider more than two equivalent Majorana degrees of freedom. In other words, for Dirac neutrinos, we have to sum the infinite number of terms, like in Eqs. (4) and (4), each of them being a branching diagram. This task seems to be quite challenging.